Teboho stands on the side of the freeway. Randall has just passed his traveling 44 m/s and he is traveling north. While Morena is headed towards him from the south at a speed of 60 m/s. Thabiso hoots his horn that has a frequency of about 400 Hz. If we consider the speed of sound in air as 340 m/s. Calculate the horn’s frequency that Randall is hearing.


1
Expert's answer
2021-05-25T14:24:48-0400

Wavelength of sound:

"\lambda=v/f=340/400=0.85\ m"

Relative speed of Morena respect to Randall is:

"\Delta v=60-44=16\ m/s"

If the horn is with Morena, then sound moving from Morena to Randall with speed:

"v_1=v+\Delta v=340+16=356\ m/s"


So, the horn’s frequency that Randall is hearing:

"f_1=v_1/\lambda=356/0.85=419\ Hz"